Diffusion Morning Report 03/25/08

Ted Berg tberg at stanford.edu
Tue Mar 25 06:11:18 PDT 2008


Good Morning All,
    Looking OK this morning. Mech pump changed on Ebeam-cryos regened, 
recovering Vacuum. TEOS 2 is having pressure control issues. LTO still 
has particle issues. Still troubleshooting Thermco 2 issues. No other 
problems at this time.

  	  	  	  	  	 
  	Date: 	3/25/08 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	  	  	  	 
Tylan 1 	up 	  	  	  	 
Tylan 2 	up 	  	  	  	 
Tylan 3 	up 	  	  	  	 
Tylan 4 	up 	  	  	  	 
Tylan 5-BBr3 	up 	  	  	  	 
Tylan 6-POCl3 	up 	  	  	  	 
Tylan 7-TEOS2 	down 	Pressure control issues 	3.71/30 	  	 
Tylan 8-SiGe 	up 	  	11.15/30 	  	 
Tylan 9-Poly 	up 	  	6.42/30 	  	 
Tylan 10-Nitride 	up 	  	6.21/30 	  	 
Tylan 11-LTO 	yellow 	particles 	6.61/30 	  	 
Tylan 12-FGA 	up 	  	  	  	 
Thermco 1 	up 	  	  	  	 
Thermco 2 	down 	-15V supply too low
	  	  	 
Thermco 3 	future 	  	  	  	 
Thermco4 	future 	  	  	  	 
Thermco Poly1 	up 	  	  	  	 
Thermco Poly2 	up 	  	  	  	 
Thermco Nitride1 	up 	  	  	  	 
Thermco Nitride2 	future 	  	  	  	 
TEL Spa 	up 	  	  	  	 
Burn Box 	Up 	  	   	  	 
Ebeam 	down 	Vacuum issues 	  	  	 
ZYGO 	up 	  	  	  	 
FGA2 	up 	  	  	  	 
SiH4 for furnaces 	up 	B on line 320 psi 	 changeover working well  	  	 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20080325/728b4e98/attachment.html>


More information about the diffusion mailing list